OpenStudy (anonymous):

Factor x2 - 36. Show your work.

OpenStudy (alexandervonhumboldt2):

ok so a^2-b^2=(a-b)*(a+b) x^2-6^2=????????

OpenStudy (anonymous):

Confused :( sorry, help ?

OpenStudy (anonymous):

6^2 = 36

OpenStudy (alexandervonhumboldt2):

ok so this is difference of squares 36=6^2 x^2-6^2= a^2-b^2=(a+b)*(a-b) substitue x for a and 6 for b

OpenStudy (anonymous):

so... a^2-6^2=(a+6)*(a-6) ...?

OpenStudy (alexandervonhumboldt2):

no a=x

OpenStudy (anonymous):

sorry I'm trying :(

OpenStudy (alexandervonhumboldt2):

substitue x instead of a

OpenStudy (anonymous):

x^2-6^2=(x+6)*(x-6) ?

OpenStudy (alexandervonhumboldt2):

YES ABSOLUTLY

OpenStudy (anonymous):

Oh yay ! So is that it ? what do I do now

OpenStudy (alexandervonhumboldt2):

x^2-36=x^2-6^2=(x+6)*(x-6)

OpenStudy (anonymous):

so it's x+6 and x-6